X³-Y³=72.XY=6 find X-Y= ? Please solve it .fast anyone can solve it?

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
(x-y)^3 = x^3 – y^3 – 3xy(x-y)
                   = 72-3(6)(x-y)
                   = 72-18(x-y)
Let (x-y) be t
          
          t^3 = 72 – 18t
Now here it is the cubic form and tehrfore exact value cannot be determined here.
Only the root approximation be detemined.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)